# Knowledge-based Consistency Testing of Large Language Models

### v2

@@ 3.1 Knowledge Graph based Test Generation
[…]
Test Generation: Given a knowledge path $\mathbb{KG}$, KonTest leverages Algorithm 1 to exhaustively generate queries relating to all possible pairs of entities present in the path. To this end, KonTest first finds the relation, $\mathbb{KG}_{rel}$, between a pair of nodes. KonTest then generates a query, ${Query}$, using $\mathbb{KG}_{rel}$ with the aid of the template shown in Figure 3 (9 in Algorithm 1). KonTest also generates the mutated query, ${Query_{M}}$, with another template (12 in Algorithm 1). These templates are dependent on the relation between the two nodes in question. For instance, Figure 3 shows the set of queries and mutated queries (i.e., $Query_{list}$ and $MutQuery_{list}$) generated from one particular path and along with the respective templates used to generate them.
@@
12:                  ${Query_{M}}\leftarrow$ Mut_Gen(${Query}$)
@@ 3.2 Test Oracles
To check the consistency in SUT’s knowledge graph, Algorithm 2 first inspects whether a direct path exists between any pair of nodes ($Node_{1},Node_{2}$) in ${Graph_{Gen}}$ (7 in Algorithm 2). In the absence of such a direct path, when an indirect path exists between the same pair of nodes (11 in Algorithm 2), KonTest detects an ontological consistency error and updates the error count, ${Err_{count}}$. Intuitively, a direct path between two nodes indicates that the SUT responded positively to a question relating the two nodes. Similarly, the presence of an indirect path indicates that SUT knows that a positive relationship exists between the nodes albeit through more than one queries. As such, the SUT can be considered to be exhibiting an ontological inconsistency. As an example, consider the ontological oracle illustrated in Table 1. In this case, if an indirect path exists between nodes for Kinawley and Ireland via node Ulster (i.e., the responses to both Q3 and Q4 are positives from the SUT), then KonTest concludes that a direct path should also exist between Kinawley and Ireland. Hence, the absence of a positive response for the respective query (Q2) is indicated as an ontological error. As a byproduct of our ontological oracle, KonTest computes the knowledge coverage ${Coverage_{count}}$ for the SUT.
@@ 3.3 Knowledge Gap Mitigation
[…]
@@ 5 Evaluation Results
LLM (Subject) Valid Test Executions (%) Errors (%)
@@
LLM (Subject) Knowledge Gap (%) Overall Coverage (%)

### v3

@@ 3.1 Knowledge Graph based Test Generation
Knowledge Graph Construction: KonTest allows the developer to specify the list of entities that are of interest. With these initial set of entities, KonTest then queries an external source of knowledge to compute, up to a certain depth, all other related entities for the considered relation. Developers can also control this depth, allowing them to determine the degree to which the vicinity of the initial set of entities is explored. Once the initial knowledge base is constructed, KonTest extracts the full path, $\mathbb{KG}$, associated with a leaf node of the knowledge base. This path is then used to guide the test generation process. For example, given the leaf entity Kinawley, as shown in Figure 1, KonTest extracts the $\mathbb{KG}$ for Kinawley as $\mathit{Kinawley}\rightarrow\mathit{Ulster}\rightarrow\mathit{Ireland}$.
Test Generation: Given a knowledge path $\mathbb{KG}$, KonTest leverages Algorithm 1 to exhaustively generate queries relating to all possible pairs of entities present in the path. To this end, KonTest first finds the relation, $\mathbb{KG}_{rel}$, between a pair of nodes. KonTest then generates a query, Query{Query}, using $\mathbb{KG}_{rel}$ with the aid of the template shown in Figure 3 (˜9 in Algorithm 1). KonTest also generates the mutated query, $Query_{M}$}, with another template (˜12 in Algorithm 1). These templates are dependent on the relation between the two nodes in question. For instance, Figure 3 shows the set of queries and mutated queries (i.e., $Query_{list}$ and $MutQuery_{list}$) generated from one particular path and along with the respective templates used to generate them.
@@
12:      $Query_{M}$$\leftarrow$ Mut_Gen(Query{Query})
@@ 3.2 Test Oracles
To check the consistency in SUT’s knowledge graph, Algorithm 2 first inspects whether a direct path exists between any pair of nodes ($Node_{1},Node_{2}$) in $Graph_{Gen}$} (˜7 in Algorithm 2). In the absence of such a direct path, when an indirect path exists between the same pair of nodes (˜11 in Algorithm 2), KonTest detects an ontological consistency error and updates the error count, $Err_{count}$}. Intuitively, a direct path between two nodes indicates that the SUT responded positively to a question relating the two nodes. Similarly, the presence of an indirect path indicates that SUT knows that a positive relationship exists between the nodes albeit through more than one queries. As such, the SUT can be considered to be exhibiting an ontological inconsistency. As an example, consider the ontological oracle illustrated in Table 1. In this case, if an indirect path exists between nodes for Kinawley and Ireland via node Ulster (i.e., the responses to both Q3 and Q4 are positives from the SUT), then KonTest concludes that a direct path should also exist between Kinawley and Ireland. Hence, the absence of a positive response for the respective query (Q2) is indicated as an ontological error. As a byproduct of our ontological oracle, KonTest computes the knowledge coverage $Coverage_{count}$} for the SUT.
@@ 3.3 Knowledge Gap Mitigation
We propose a weighted ensemble method that leverages the test suite generated by KonTest to mitigate the discovered knowledge gaps. The ensemble method uses the number of metamorphic errors found for each SUT on a per relation basis to inform its scoring system. A relation that induces $x$ errors is given a score of 5x5-x since each relation can maximally induce five errors. The cumulative SUT specific score, $\mathit{Score}_{i}$, is then found by summing the scores for each relation. Next, the relative weight, $\mathit{W}_{i}$, for each SUT in the set of SUTs ($\mathbb{SUT}$), is computed via the following equation:
$\mathit{W}_{i}=\frac{\mathit{Score}_{i}}{\sum\mathit{Score}_{i}}\forall i\in\mathbb{SUT}$ (1)
@@ 5 Evaluation Results
LLM
Valid Test Executions (%) Errors (%)
@@
LLM
Knowledge Gap (%)
Overall
Coverage (%)
@@ 9 Ethics Statement
10 Acknowledgment
This work was partially funded by grant number SMU-SUTD 2023_02_04 and the Singapore Ministry of Education (MOE) Present’s Graduate Fellowship. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the view of the respective funding agencies.